1/9+1/2 divided by 1/6 x 5/12
Answer by stanbon(75887)

1/9+1/2 divided by 1/6 x 5/12
------
[(1/9)+(1/2)] / [(1/6)(5/12)]
--------
[(2*1+9*1)/18] / (5/72)
--------
= (11/18)*(72/5)
--------
= (11/5)*4
-----
= 44/5
---
= 8.8
----------------
Cheers,
Stan H.